The number 50 is divided into two equal parts, the ratio of the largest part to the smallest is 2/3, then the largest part is?
1
Expert's answer
2011-12-27T08:10:11-0500
First of all, in the question it's said that "number 50 is divided into two equal parts" and though "ratio of the largest part to the smallest is 2/3". obviously, two equal parts are equal hence they are 25,25. If we drop it off and consider the second part saying about ratio then we will act in the following way. Let's take a look.

Let's denote L one part and m another one. Then

s/L = 2/3
L+s = 50 ==> s = 50-L

(50-L)/L = 2/3

50/L - 1 = 2/3

50/L = 5/3

L/10 = 3

L = 30

s = 50-30 = 20.

So, the large part is 30 and the small part is 20.